# RePO: ReLU-based Preference Optimization

### v1

@@ 6 Discussion
Conclusion. In this work, we propose RePO, a simple preference optimization by replacing logistic weighting with a ReLU-based max-margin loss, eliminating the need for hyperparameter $\beta$. Theoretically, RePO is characterized as SimPO’s $\beta\rightarrow\infty$ limit, collapsing to binary thresholding for data filtering. Empirically, it matches or outperforms DPO and SimPO on AlpacaEval 2 and Arena-Hard with only a single tunable parameter ($\gamma$). Its gradient dynamics prioritize challenging samples, mitigating over-optimization while maintaining training stability. This work establishes RePO as a robust, hyperparameter-efficient alternative for offline alignment.
Limitations and Future Work. One limitation of our approach is that it currently operates in an offline setting, and future research will focus on extending it to online reinforcement learning frameworks. Integrating repo with online RL techniques could enhance real-time adaptability and scalability. Another direction for future work is to explore how the cutoff point strategy can be further refined to sustain performance improvements in self-play scenarios (Chen et al., 2024b), particularly in highly dynamic environments.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Related Works
Reinforcement learning from human feedback. RLHF is a technique designed to align large language models with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Azar et al., 2023). Traditional RLHF is typically structured in three stages: supervised fine-tuning (Zhou et al., 2023; Taori et al., 2023; Geng et al., 2023; Conover et al., 2023; Köpf et al., 2023; Ding et al., 2023), reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024a; Lightman et al., 2023; Havrilla et al., 2024; Lambert et al., 2024), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). In the third stage, Proximal Policy Optimization (PPO) is widely adopted. In contrast, RLOO (Ahmadian et al., 2024) reduces the GPU memory footprint of RLHF by eliminating the Critic model and leverages a Leave-One-Out strategy to achieve superior performance. GRPO (Shao et al., 2024), another variant of PPO, improves mathematical reasoning abilities while optimizing memory usage by replacing the Leave-One-Out method with a direct subtraction of the mean of all samples for a given prompt.
@@
Offline preference optimization. In addition to DPO, several alternative preference optimization objectives have been proposed. IPO (Azar et al., 2023) addresses overfitting issues inherent in DPO. ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) aim to remove reliance on a reference model. R-DPO (Park et al., 2024) targets the reduction of exploitation due to sequence length, while KTO (Ethayarajh et al., 2024) handles preference optimization in the absence of pairwise data. CPO (Xu et al., 2024) and $\beta$-DPO (Wu et al., 2024b) focus on improving the quality of preference data. Another research direction addresses noise in offline alignment, which arises from the need to construct pairwise data. rDPO (Chowdhury et al., 2024), a variant of DPO, mitigates preference noise and enhances policy robustness, while DrDPO (Wu et al., 2024a) applies distributed robust optimization to tackle this issue. Other works have approached the problem through divergence regularization (Wang et al., 2024a; Huang et al., 2024), selection of high-quality data (Khaki et al., 2024; Ye et al., 2024), or reweighting loss functions (Wang et al., 2024c; Muldrew et al., 2024; Yang et al., 2024).
@@
Iterative Preference Optimization. Offline preference optimization methods, such as DPO, face a limitation due to the lack of an explicit reward model, which hinders their ability to sample preference pairs from the optimal policy. To address this, iterative preference optimization techniques have been proposed. These methods iteratively update the reference model using the most recent policy model or generate new preference pairs in each iteration (Dong et al., 2024; Kim et al., 2024; Rosset et al., 2024; Xiong et al., 2024; Yuan et al., 2024; Chen et al., 2024b; Wu et al., 2024c; Gao et al., 2024). For instance, SPIN (Chen et al., 2024b) employs a self-play framework to fine-tune the model in a supervised manner, while Yuan et al. (2024) annotate preferences throughout the iterative process. REBEL improves sample quality by regressing the relative reward. Additionally, (Calandriello et al., 2024) generates data using a mixture policy, similar to the Nash-MD algorithm (Rosset et al., 2024).
Appendix B Proofs
B.1 Proof of Lemma 3.1
@@ Appendix B Proofs
Here, $-M_{\theta}+\gamma>0$. As $\beta\to\infty$,
@@
Here, $-M_{\theta}+\gamma<0$. As $\beta\to\infty$,
@@
[…]
@@
3. Epigraph Characterization: The epigraph $\mathrm{epi}(h)$ coincides with the convex hull of $\mathrm{epi}(\mathcal{L}_{0\text{-}1})\cap(D\times\mathbb{R})$. The affine segment $h(x)=-\frac{x}{a}$ on $[-a,0)$ connects the points $(-a,1)$ and $(0,0)$, forming the tightest convex fit to the 0-1 loss’s discontinuity. By Theorem 1 in (Princeton University, Accessed 2025), this construction achieves the convex envelope. ∎
B.3 Proof of Corollary 4.3
@@
Part 2: For $D=[-a,b]$, both $\mathcal{L}_{0\text{-}1}(x)$ and $\frac{1}{a}\mathrm{ReLU}(-x)$ attain their minimum value 0 on $[0,b]$. For $x\in[-a,0)$, $\frac{1}{a}\mathrm{ReLU}(-x)$ is strictly decreasing, achieving its minimum at $x=0$. Thus:
@@
We demonstrate violation of the convex envelope’s defining property. Consider $D=[-1,1]$:
@@ 1. Underestimation Failure: For $x>0$:
$-\log\sigma(x)=-\log\left(\frac{1}{1+e^{-x}}\right)=\log(1+e^{-x})>0=\mathcal{L}_{0\text{-}1}(x)$
@@
Appendix C Experiments
C.1 Implementation Details
@@
  • Optimizer: Adam (Kingma & Ba, 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
@@
DPO (Rafailov et al., 2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in[0.01,0.05,0.1]$
IPO (Azar et al., 2023) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
CPO (Xu et al., 2024) $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
KTO (Ethayarajh et al., 2024) $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$
@@
ORPO (Hong et al., 2024) $-\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$ $\lambda\in[0.1,0.5,1.0,2.0]$
@@
R-DPO (Park et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
@@
SimPO (Meng et al., 2024) $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in[2.0,4.0,6.0,8.0]$
@@
RePO $\text{ReLU}[-(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma)]$ $\gamma\in[0.2,0.4,0.6,0.8]$
@@
C.2 Downstream Task Evaluation
@@
To assess the impact of RePO on downstream task performance, we evaluate models trained with different preference optimization methods on a diverse set of tasks from the Huggingface Open Leaderboard (Beeching et al., 2023). The tasks include MMLU (Hendrycks et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2022), Winograd (Levesque et al., 2012), and GSM8K (Cobbe et al., 2021). We adhere to standard evaluation protocols and present the results for all models in Table 5.
@@
Conclusion and Future Work. In summary, RePO shows promising results across a range of downstream tasks, with notable improvements in tasks related to commonsense reasoning and truthfulness. However, it faces challenges in preserving general knowledge and mathematical reasoning. Future work could focus on incorporating additional strategies to mitigate performance drops in these areas, possibly through the integration of task-specific objectives or enhanced fine-tuning techniques. Additionally, further empirical studies are needed to explore the impact of different preference optimization datasets and objectives on downstream performance.
C.3 RePO with varying $\gamma$

### v2

@@ 6 Discussion
Conclusion Our exploration of simple ReLU activation in preference learning has revealed several key insights. We found that binary thresholding, implemented through a straightforward ReLU function, provides an effective mechanism for preventing over-optimization in language model alignment. Our theoretical analysis showed that this seemingly simple approach is, in fact, the convex envelope of the ideal 0-1 loss function, explaining its surprising effectiveness. Rather than developing yet another complex preference optimization method, our work uncovered how fundamental properties like data selection and implicit curriculum learning emerge naturally from basic principles.
Limitations and future directions. Our current exploration is limited to offline preference learning settings. Future work could investigate how these insights might extend to online learning scenarios, where preferences are gathered interactively. Additionally, while we found that a fixed margin threshold works well in practice, exploring adaptive or context-aware thresholds might further improve performance in highly dynamic environments. The relationship between binary filtering and self-play scenarios [27] — where the model generates its own feedback — is another promising direction that could lead to more scalable alignment techniques.
Beyond alignment, our work connects to LLM reasoning research [28, 29]. Future work should investigate how KL penalties and gradient clipping in GRPO [30] and PPO [5] balance preventing over-optimization against preserving reasoning capabilities — a critical consideration for advancing alignment methodologies.
Acknowledgments and Disclosure of Funding
This research was supported by the National Science and Technology Major Project (2023ZD0121102), the National Natural Science Foundation of China (U24B20180, 62302321), and the Fundamental Research Funds for the Central Universities (WK2100250065). This research also benefited from the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Appendix A Related Works
Reinforcement learning from human feedback. RLHF is a technique designed to align large language models with human preferences and values [31, 32, 33, 20]. Traditional RLHF is typically structured in three stages: supervised fine-tuning [34, 35, 36, 37, 38, 39], reward modeling [9, 40, 41, 42, 43, 44], and policy optimization [5, 45]. In the third stage, Proximal Policy Optimization (PPO) is widely adopted. In contrast, RLOO [46] reduces the GPU memory footprint of RLHF by eliminating the Critic model and leverages a Leave-One-Out strategy to achieve superior performance. GRPO [47], another variant of PPO, improves mathematical reasoning abilities while optimizing memory usage by replacing the Leave-One-Out method with a direct subtraction of the mean of all samples for a given prompt.
@@
Offline preference optimization. In addition to DPO, several alternative preference optimization objectives have been proposed. IPO [20] addresses overfitting issues inherent in DPO. ORPO [23] and SimPO [11] aim to remove reliance on a reference model. R-DPO [24] targets the reduction of exploitation due to sequence length, while KTO [22] handles preference optimization in the absence of pairwise data. CPO [21] and $\beta$-DPO [48] focus on improving the quality of preference data. Another research direction addresses noise in offline alignment, which arises from the need to construct pairwise data. rDPO [49], a variant of DPO, mitigates preference noise and enhances policy robustness, while DrDPO [50] applies distributed robust optimization to tackle this issue. Other works have approached the problem through divergence regularization [51, 52], selection of high-quality data [53, 54], or reweighting loss functions [55, 56, 57].
@@
Iterative Preference Optimization. Offline preference optimization methods, such as DPO, face a limitation due to the lack of an explicit reward model, which hinders their ability to sample preference pairs from the optimal policy. To address this, iterative preference optimization techniques have been proposed. These methods iteratively update the reference model using the most recent policy model or generate new preference pairs in each iteration [58, 59, 60, 61, 62, 27, 63, 64]. For instance, SPIN [27] employs a self-play framework to fine-tune the model in a supervised manner, while Yuan et al. [62] annotate preferences throughout the iterative process. REBEL improves sample quality by regressing the relative reward. Additionally, [65] generates data using a mixture policy, similar to the Nash-MD algorithm [60].
Appendix B Broader Impacts
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here
Appendix C Proofs
C.1 Proof of Lemma 3.1
@@ Appendix C Proofs
Here, $M_{\theta}+\gamma>0$. As $\beta\to\infty$,
@@
Here, $M_{\theta}+\gamma<0$. As $\beta\to\infty$,
@@
$=-\mathbb{E}_{\mathcal{D}}\left[\lim_{\beta\to\infty}\beta s_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]\quad\text{(Dominated Convergence Theorem)}$ (23)
@@
3. Epigraph Characterization: The epigraph $\mathrm{epi}(h)$ coincides with the convex hull of $\mathrm{epi}(\mathcal{L}_{0\text{-}1})\cap(D\times\mathbb{R})$. The affine segment $h(x)=-\frac{x}{a}$ on [a,0)[-a,0) connects the points (a,1)(-a,1) and $(0,0)$, forming the tightest convex fit to the 0-1 loss’s discontinuity. By Theorem 1 in [66], this construction achieves the convex envelope. ∎
C.3 Proof of Corollary 4.3
@@
Part 2: For D=[a,b]D=[-a,b], both $\mathcal{L}_{0\text{-}1}(x)$ and $\frac{1}{a}\mathrm{ReLU}(-x)$ attain their minimum value 0 on $[0,b]$. For $x\in[-a,0)$, $\frac{1}{a}\mathrm{ReLU}(-x)$ is strictly decreasing, achieving its minimum at $x=0$. Thus:
@@
We demonstrate violation of the convex envelope’s defining property. Consider D=[1,1]D=[-1,1]:
@@ 1. Underestimation Failure: For $x>0$:
$\log\sigma(x)=-\log\left(\frac{1}{1+e^{-x}}\right)=\log(1+e^{-x})>0=\mathcal{L}_{0\text{-}1}(x)$
@@
Appendix D Experiments
D.1 Implementation Details
@@
  • Optimizer: Adam [13] ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
@@
SLiC-HF [7] $\max\left(0,\delta-\log\pi_{\theta}(y_{w}|x)+\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda\in[0.1,0.5,1.0,10.0]$
$\delta\in[0.1,0.5,1.0,2.0]$
DPO [6] $\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in[0.01,0.05,0.1]$
IPO [20] $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
CPO [21] $\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
KTO [22] $\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$
@@
ORPO [23] $\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$ $\lambda\in[0.1,0.5,1.0,2.0]$
@@
R-DPO [24] $\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
@@
SimPO [11] $\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in[2.0,4.0,6.0,8.0]$
@@
RePO $\text{ReLU}[-(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma)]$ $\gamma\in[0.2,0.4,0.5,0.6,0.8]$
@@
D.2 Downstream Task Evaluation
@@
To assess the impact of RePO on downstream task performance, we evaluate models trained with different preference optimization methods on a diverse set of tasks from the Huggingface Open Leaderboard [25]. The tasks include MMLU [67], ARC [68], HellaSwag [69], TruthfulQA [70], Winograd [71], and GSM8K [72]. We adhere to standard evaluation protocols and present the results for all models in Table˜5.
@@
D.3 RePO with varying $\gamma$